OpenStudy (agentnao):

What is the 32nd term of the arithmetic sequence where a1 = –33 and a9 = –121?

OpenStudy (anonymous):

i think -374

OpenStudy (angela210793):

a9=a1+(n-1)*d -121=-33+8d d=-11 a32=a1+31*(-11)=-374 Math is right :)

OpenStudy (angela210793):

an=a1+(n-1)*d

OpenStudy (agentnao):

Oh wow. I feel like such an idiot! I know what I did wrong now. Thank you so much :)

OpenStudy (angela210793):

yw :D

OpenStudy (anonymous):

a32=-374